OpenStudy (anonymous):

A scale on a map shows that 2 inches = 25 miles. Part A How many inches on the map represent 60 miles? Enter your answer in the box. inches Part B How many miles are represented by 1 4 inch on the map? Enter your answer in the box. miles

OpenStudy (anonymous):

@Directrix @Vocaloid

OpenStudy (anonymous):

@Mehek14

Mehek (mehek14):

\(\tt{\dfrac{25}{2}=\dfrac{60}{x}}\) cross multiply

OpenStudy (anonymous):

i got 30 for 2 and 60 but i dont know what to do for 25 and x

OpenStudy (anonymous):

@Mehek14

Mehek (mehek14):

for A 60 * 2 = 120 120/25 = ?

OpenStudy (anonymous):

4.8

Mehek (mehek14):

so A is 4.8

OpenStudy (anonymous):

okay and b?

Mehek (mehek14):

for B i think 50 is correct